OpenStudy (anonymous):

Given that 2^(2x+2)⋅5^(x−1)=8^x⋅5^2x , evaluate 10 power x.

terenzreignz (terenzreignz):

\[\large 2^{2x}+2\cdot5^{x-1}=8^{x}\cdot5^{2x}\]

OpenStudy (anonymous):

2^2x+2

OpenStudy (anonymous):

x 5^x-1

OpenStudy (anonymous):

= 8 power x X 5 power 2x

terenzreignz (terenzreignz):

\[\large 2^{2x+2}+2\cdot5^{x-1}=8^{x}\cdot5^{2x}\]

OpenStudy (anonymous):

it's not +2 it's multiply

OpenStudy (anonymous):

without a 2

terenzreignz (terenzreignz):

\[\large 2^{2x}\cdot5^{x-1}=8^{x}\cdot5^{2x}\]

OpenStudy (anonymous):

wrong!

OpenStudy (anonymous):

2^2x+2 Multiply 5 power x-1 = ...

terenzreignz (terenzreignz):

\[\large 2^{2x}+2\cdot5^{x-1}=8^{x}\cdot5^{2x}\] Back here, then?

OpenStudy (anonymous):

2 power (2x+2) x 5 power (x-1) = 8 power x X 5^2x

terenzreignz (terenzreignz):

\[\large 2^{2x+2}\cdot5^{x-1}=8^{x}\cdot5^{2x}\]

OpenStudy (anonymous):

finally :P

terenzreignz (terenzreignz):

Sorry, you know, it's a bit hard to understand when written plainly... now hang on, let me see if this is something I'm capable of :D

OpenStudy (anonymous):

remember to evaluate 10 power x

terenzreignz (terenzreignz):

Something tells me this isn't solvable...

OpenStudy (anonymous):

i have the answer.

OpenStudy (anonymous):

it's not asking you to solve , evaluate 10 to the power of x

terenzreignz (terenzreignz):

\[\large \frac{2^{x+2}\cdot2^{x}\cdot5^{x}}{5}=8^{x}\cdot5^{2x}\]

terenzreignz (terenzreignz):

\[\large \frac{2^{x+2}\cdot10^{x}}{5}=8^{x}\cdot5^{2x}\]

terenzreignz (terenzreignz):

\[\large 10^x=2^{3x}\cdot5^{2x}\cdot5\cdot2^{-x-2}\]

terenzreignz (terenzreignz):

And then evaluate, I reckon. I don't think there's a value for x that would satisfy this anyway... I'm so lost :D

OpenStudy (anonymous):

then how?

OpenStudy (anonymous):

answer is 4/5

OpenStudy (anonymous):

2 power x multiply 5 power x = 4/5

terenzreignz (terenzreignz):

I don't think that holds... I'm missing something, probably

OpenStudy (anonymous):

\[2^{2x + 2} \times 5^{x-1} = 8^{x} \times 5^{2x}\]

OpenStudy (anonymous):

the 8 power x would become 2^3x

terenzreignz (terenzreignz):

and it should become 3x = 2x + 2 x = 2

OpenStudy (anonymous):

you cant assume that way x is not equal to 2 i did that and it was a mistake.

terenzreignz (terenzreignz):

But I tried x = 4/5, it didn't work, neither, try it in your calculator.

OpenStudy (anonymous):

dude 10 to the power of x = 4/5

terenzreignz (terenzreignz):

oh, so it's 10^x then... hang on...

terenzreignz (terenzreignz):

Right, I honestly have no idea how to go about this :/ I am shamed :(

terenzreignz (terenzreignz):

Does this involve logarithms, by any chance?

terenzreignz (terenzreignz):

Ok, never mind, I've reached clarity now :D

terenzreignz (terenzreignz):

\[\large 10^x=2^{3x}\cdot5^{2x}\cdot5\cdot2^{-x-2}\]\[\large 10^x=2^{2x}\cdot2^{x}\cdot5^{2x}\cdot5\cdot2^{-x-2}\]\[\large 10^x=2^{2x}\cdot5^{2x}\cdot2^{x}\cdot5\cdot2^{-x-2}\]\[\large 10^x=10^{2x}\cdot5\cdot2^{x-x-2}\]\[\large 10^x=10^{2x}\cdot5\cdot2^{-2}\]

terenzreignz (terenzreignz):

And now, it's easier to see that 10^x = 4/5 It turns out I gave up too soon, better not do that again :D

OpenStudy (anonymous):

oh i see...

terenzreignz (terenzreignz):

\[\large 10^x=\frac{5}{4}\cdot10^{2x}\]\[\large 0=\frac{5}{4}\cdot10^{2x}-10^x\]\[\large 0=10^x\left(\frac{5}{4}\cdot10^{x}-1\right)\]

terenzreignz (terenzreignz):

So either \[\huge 10^x = 0 \ \ or \ \ 10^x = \left(\frac{5}{4}\right)^{-1}\] And it can't be the former.

OpenStudy (anonymous):

Thank you!

terenzreignz (terenzreignz):

No problem